OpenStudy (anonymous):

2x+3y=5 help me find solutions to graph. #MEDAL WILL BE REWARDED!!!!!

OpenStudy (anonymous):

2x+3y=5 2x/5 + 3y/5=1 in intercept form: \[x/(5/2) + y/(5/3) = 1\] hence draw a straight line joining points: (5/2,0) and (0,5/3) the line that you get after joining the two points is the desired line

OpenStudy (anonymous):

To find x-intercept just plug y = 0 in equation & solve for x & to find y -intercept plug x=0 & solve for y , So you will get intercepts like x intercept (5/2,0) & y-intercept (0 , 5/3)

OpenStudy (anonymous):

& then plot & join these points by line

OpenStudy (anonymous):

ok but how would u gragh 5/2. like what time of number scale

OpenStudy (anonymous):

Just divide 5 by 2 , what do you get ?

OpenStudy (anonymous):

2.5 but 5/3 is 1.667

OpenStudy (anonymous):

Yes correct !! , so we can plot this points right ?

OpenStudy (anonymous):

yes ,you cant draw an exact graph, just an approx by this method

OpenStudy (anonymous):

no becuz how would u organize the number scale

OpenStudy (anonymous):

0.5 , 1 , 1.5 , 2.5 like this Or we can round off 1.6667 to 1.7

OpenStudy (anonymous):

# @tgawade Thnx

OpenStudy (anonymous):

You are welcome :)

OpenStudy (anonymous):

medal! medal!

OpenStudy (anonymous):

srry i awardes to best answer already @hanzi

OpenStudy (anonymous):

its okay , i can bear the wrath ;)

OpenStudy (anonymous):

i got my questions im posting on rational equations